OpenStudy (anonymous):

find the limit as x approaches 0 for ((x^2)(cotx))/sinx

OpenStudy (anonymous):

((x^2)(cotx))/sinx= (x*cotx)*(x/sinx)= (x*cosx/sinx)*(x/sinx)=cosx*(x/sinx)^2 now as x approaches 0 cosx and x/sinx approaches 1 hence required limit is 1*1^2=1

OpenStudy (anonymous):

I agree with matricked.